OpenStudy (anonymous):

Fan&Medal Two executives fly towards each other for a meeting. The first jet goes 320 mph, while the second goes 375 mph. If they started from opposite coasts 3100 miles apart, how long will it take them to meet? Round to the nearest tenth if necessary. Select one: a. 9.7 hours b. 8.3 hours c. 2.2 hours d. 4.5 hours

OpenStudy (anonymous):

d

OpenStudy (anonymous):

do you want to know why

OpenStudy (anonymous):

surec:

OpenStudy (anonymous):

Let the first jet cover x miles in t hours Then in same time the second jet will have covered 3100-x miles in same time t hours Now time = distance/speed time for first jet t = x/320 ---------------------- (I) time for second jet t = (3100-x)/375 -------------- (II) since t is same for both, we can equate the two x 3100-x ----- = --------- 320 375 Cross-multiplying 375x = 320(3100-x) 375x = 992000 - 320x 375x + 320x = 992000 695x = 992000 x= 992000/695 x = 1427.338 substituting in (I) t = 1427.338/320 t= 4.46 so ans is 4.5 hours

OpenStudy (anonymous):

woaa. Thanks man:3

OpenStudy (anonymous):

your welcome